Given

[1 1 -1]

A=[2 2 3]

[4 0 0]

[1 1 -1]

b=[2 2 3]

[4 0 0]

1.Find –A-1+ 3BT

2.Find B-1+ ( AT+A-1)


1
Expert's answer
2022-03-19T02:41:29-0400


Given that


"A=\begin{bmatrix}
 1 & 1 & -1\\
 2 & 2 & 3\\
 4 & 0 &0
\end{bmatrix}" and "B=\begin{bmatrix}
 1 & 1 & -1\\
 2 & 2 & 3\\
 4 & 0 &0
\end{bmatrix}"


Then


"A^{T}=\begin{bmatrix}
 1 & 2 & 4\\
 1 & 2 & 0\\
 -1 & 3 &0
\end{bmatrix}" and "B^{T}=\begin{bmatrix}
 1 & 2 & 4\\
 1 & 2 & 0\\
 -1 & 3 &0
\end{bmatrix}"


"det(A)=\begin{vmatrix}
1 & 1 & -1\\
 2 & 2 & 3\\
 4 & 0 &0
\end{vmatrix}"


"det(A)=(1)(0-0)-(1)(0-12)+(-1)(0-8)"


"det(A)=0+12+8"


"det(A)=20"


And


"det(B)=\begin{vmatrix}
1 & 1 & -1\\
 2 & 2 & 3\\
 4 & 0 &0
\end{vmatrix}"


"det(B)=(1)(0-0)-(1)(0-12)+(-1)(0-8)"


"det(B)=0+12+8"


"det(B)=20"




For the matrix "A=\begin{bmatrix}
 1 &1&-1 \\
 2 & 2&3\\
4&0&0
\end{bmatrix}"


"A^{-1}=\frac{adjointA}{det(A)}"


The adjoint of A is calculated by the cofactor method


"A^{-1}=\frac{\begin{bmatrix}
 0 &0&5 \\
 12&4&-5 \\
-8 & 4&0
\end{bmatrix}}{(20)}"


"A^{-1}=\begin{bmatrix}
 \frac{0}{20} &\frac{0}{20}&\frac{5}{20} \\
 \frac{12}{20} & \frac{4}{20}&\frac{3}{20}\\
\frac{-8}{20}&\frac{4}{20}&\frac{0}{20}
\end{bmatrix}"


"A^{-1}=\begin{bmatrix} 0 &0 & 0.25\\ 0.6 & 0.2 & -0.25\\ -0.4 & 0.2 &0 \end{bmatrix}"


Similarly


for the matrix "B=\begin{bmatrix}
 1 &1&-1 \\
 2 & 2&3\\
4&0&0
\end{bmatrix}"


"B^{-1}=\frac{adjointB}{det(B)}"


The adjoint of B is calculated by the cofactor method


"B^{-1}=\frac{\begin{bmatrix}
 0 &0&5 \\
 12&4&-5 \\
-8 & 4&0
\end{bmatrix}}{(20)}"


"B^{-1}=\begin{bmatrix}
 \frac{0}{20} &\frac{0}{20}&\frac{5}{20} \\
 \frac{12}{20} & \frac{4}{20}&\frac{3}{20}\\
\frac{-8}{20}&\frac{4}{20}&\frac{0}{20}
\end{bmatrix}"


"B^{-1}=\begin{bmatrix} 0 &0 & 0.25\\ 0.6 & 0.2 & -0.25\\ -0.4 & 0.2 &0 \end{bmatrix}"


Solution (A)



Solution (B)